### v2

Abstract
Disaster support mobile apps play a central role during emergencies in conveying timely alerts to communities and providing recovery guidance. Their importance continues to grow as climate change increases the frequency and intensity of disasters, and as mobile devices become a primary information source for most people. However, despite their critical role, there has been limited systematic investigation into how these apps are designed and how users experience them in real-world contexts. This paper presents a large-scale empirical study of disaster support mobile apps aimed at addressing this gap. We investigated 44 apps available on the Google Play Store and the App Store and analyzed 28,161 user reviews to understand common functionalities, user expectations, and challenges encountered during use. Our feature analysis identifies 13 core capabilities spanning the stages of preparation, response, and recovery of the disaster lifecycle. Topic Modelling revealed 18 topics, with the highest discussions focusing on app alert functionality, the use of maps, and monetary transactions. Sentiment analysis revealed that while 53.09% of users provided positive feedback, 24.30% expressed negative opinions, and 11.25% remained neutral. It also showed that signup/signin issues, network problems, and app crashes were the most frustrating to users. These impacted user safety by preventing access when it mattered most. Our findings emphasize the need to approach disaster apps as safety-critical and human-centred systems, where reliability and ease of use are essential. We provide practical recommendations to support developers and emergency agencies in improving the dependability and accessibility of future disaster support apps.
keywords:
disaster support mobile apps, safety-critical software, feature analysis, human-centered software
1 Introduction
Natural disasters, including floods, heat waves, earthquakes, and wildfires, have become increasingly common. According to the Atlas of Mortality and Economic Losses from Weather [UNFCCC2021], natural disasters have increased fivefold over that 50-year period. Such disasters lead to loss of life and livelihood, disrupt education, harm health and hinder economic growth.
However, many communities can reduce the impact of these disasters by improving their resilience through better preparation and access to information on disaster management strategies before, during, and after disasters. Traditionally, such information was shared through broadcast media such as television, radio, or word of mouth, and later through short message service (SMS) alerts. While SMS can deliver brief warnings, mobile apps now provide more advanced, interactive, and adaptable platforms. With 85% of today’s global population using mobile phones, disaster-related information is increasingly delivered through mobile apps [HowManyP98_online, tan2020usability].
Disaster applications are increasingly regarded as critical support tools due to their potential to influence life or death situations [tan2020understanding, dallo2021should]. Consequently, many individual developers, emergency agencies, and government organizations have begun developing and publishing their own versions of disaster apps on the Google Play Store and the App Store. However, several issues have been reported in these apps, such as complicated interfaces, missed alerts and overwhelming information load [CFAappgl73_online, tan2020understanding, tan2020usability]. In acute situations such as disaster-related emergencies, even seemingly minor issues can become critical concerns that can lead to a decision between life and death.
Consider an example of a poorly designed disaster app. Firstly, due to bugs in the app, it may issue delayed bush fire alerts or completely miss alerts, which can lead to loss of lives and livelihoods [TheGuardian2025]. Secondly, poorly designed map visualizations, such as rainbow color palettes, can mislead users. Studies show that non-uniform color scales in flood hazard maps distort how people judge severity and extent of risks [Jenny2024]. Thirdly, due to privacy and security issues, the app may get hacked which can lead to issuing incorrect alerts that leads to massive community panics [Heckel2020]. Therefore, it is important to develop these apps methodically and by following good software engineering principles, and to ensure that key disaster support features are built to a high standard.
When developing these apps, rather than re-inventing the wheel and each local emergency agency developing a disaster app from scratch, they can benefit from referring to existing apps. Moreover, since disasters often share similar patterns across regions, the findings from one country can inform improvements in others. However, to the best of authors’ knowledge, a study analyzing the wide range of disaster app features and their reviews on both the Google Play Store and the App Store is currently lacking. In this study we aim to address this research gap.
We present a detailed analysis of disaster-related mobile apps available on the Google Play Store and the App Store, and provide a comprehensive list of commonly implemented features across these apps. This work aims to offer emergency agencies a feature pool that can guide the development of their own applications. In addition, we analyze user reviews using topic modeling techniques to identify the key themes discussed by users. A sentiment analysis is then performed to uncover common preferences and pain points, providing insights into the aspects of these apps that require greater attention.
The rest of this paper is structured as follows: Section 2 describes related work, Section 3 outlines the research design and Section 4 presents the results of our analysis. Section 5 presents an overall discussion and implications of our findings in Section 6, threats to validity and in Section 7, a summary with key future directions.
2 Related Work
2.1 Mobile Apps for Disaster Management
Disaster apps have a significant diversity in the type of information they provide and their intended target groups. The type of information vary based on type of disaster and the stage of disaster lifecycle they focus on [dallo2021should]. The type of disaster can be a single disaster such as floods or may present information about multiple disasters in which case they are known as multi-hazard apps. The stages of a disaster lifecycle are fourfold according to disaster management literature. These are ’Preparation’, ’Response’, ’Recovery’ and ’Mitigation’ [04Unit4d19_online]. The disaster apps can focus on one or many of these stages. The preparation stage involves information such as capacity building, monitoring and providing early warnings; the response stage involves information for evacuation, search and rescue; the recovery stage involves damage assessment and reconstruction; and lastly the mitigation stage involves information such as risk analysis and risk appraisals [poser2010volunteered]. The target groups can be the global community for international apps and local communities such as state, regional communities for local apps.
Disaster apps are broadly classified into two categories: general purpose apps and purpose-built apps. General purpose apps include three sub-categories One-to-one, One-to-many and Many-to-many [Tan2017]. One-to-one refers to messaging apps such as Whatsapp and Wechat which are used by a person to send personal message to another. One-to-many refers to news apps are maintained by news agencies such as CNN and while publishing other news, they also publish disaster news to the public. Many-to-many include social media apps such as Twitter and Facebook. Out of these, people tend to favor familiar platforms such as social media which they have frequently used before the disaster occurrence [HADDOW201453].
However, disaster management authorities have concerns in promoting the use of general-purpose platforms for emergency situations, as many issues arise such as privacy, information quantity, and content quality [Tan2017]. purpose-built apps are usually maintained by emergency agencies and will focus on functions such as Alert and information dissemination: disseminating authorized information before and during disasters; and Information collation: gathering, filtering and analyzing data to build situation awareness. These apps are usually one-way; commonly originating from the authorities to the public. Examples of such apps are the NINA-Die Warn-App des BBK and the Federal Emergency Management Agency (FEMA) App, both of which were developed by their respective agencies.
2.2 App Review Analysis
Given that app reviews include rich information, a growing number of researchers and practitioners have investigated app reviews to extract actionable insights for different software engineering tasks [dkabrowski2022analysing, genc2017systematic]. App reviews have been mainly analyzed to support requirements engineering tasks. For example, one study [Tan2020Modified] collected and thematically analyzed online reviews of disaster apps. From these insights, the authors built a conceptual usability framework specific to disaster apps. Researchers [johann2017safe, de2021re, Motger2024] developed approaches to identify new features requested by users in mobile app reviews. Some have extracted non-functional requirements from app reviews (e.g., [jha2019mining, lu2017automatic]). Other work has focused on detecting key features to determine which aspects users value most or complain about the most [Wu2021]. Additional efforts have aimed at identifying issues (e.g., UI issues, crashes, bugs) that users face when working with apps by mining user reviews (e.g., [man2016experience, chen2021should, maalej2015bug, mcilroy2016analyzing, Motger2024]). Another line of research attempted to (automatically) detect fake reviews and non-informative reviews (e.g., [martens2019towards, noei2019too]). Another important research theme is the sentiment analysis of app reviews to detect users’ emotions about the app (e.g., [martens2017emotion, luiz2018feature]). Finally, some research has recently focused on identifying and classifying human, social, and ethical factors (e.g., human and ethical values [shahin2023study, obie2023automated, shams2020society, bowie2022user] and privacy [nema2022analyzing]) from app reviews.
2.3 Review-based Disaster Mobile App Analysis
Researchers have explored disaster mobile apps in a limited number of studies. Tan et al. [Tan2017] systematically reviewed crisis informatics literature, including disaster apps. However, these studies mainly focused on apps discussed in academic literature, and as we found in our analysis, not all of them were available to the public. Some were still at the proof-of-concept stage and some can’t be found in any website and application store. Corcuera et al. [corcuera_2022] also looked into disaster apps. Still, their literature review was limited to apps in the Google Play Store, and they sought only to explore the functionalities of these apps and evaluate their effectiveness for disaster risk reduction [corcuera_2022]. They did not explore the user reviews that are available in these apps for insights into user perception of features and app performance.
Studies analyzing app reviews, described in Section 2.2, typically do not cover disaster apps. This is mainly because these studies choose the reviews from the most popular apps (e.g., Instagram and TikTok apps) and app categories (e.g., health, social, communication) or randomly select reviews in the Google Play Store and the App Store. The popularity of an app is usually defined by the number of downloads, which excludes disaster apps from consideration. This exclusion also applies to other less frequently downloaded but important app categories, such as depression apps [bowie2022user] and sports apps [Chembakottu2023]. However, as explained earlier, disaster apps are used in critical situations where even a minor issue can make a difference between life and death. Therefore, a detailed study of disaster apps available to the public today is long overdue.
3 Research Design
In this study we aim to answer three key research questions. We provide these questions and our rationale for choosing them below.
3.1 Research Questions
RQ1. What are the key characteristics of disaster apps?
Rationale. We want to systematically identify, examine, and evaluate the features of disaster apps. This investigation involves a detailed review of the app functionalities, the types of natural disasters they address, and their geographic coverage. Subsequently, we analyze the apps’ popularity and quality based on metrics such as the number of downloads and user ratings and explore their monetization strategies. This analysis will help emergency agencies and their app developers to understand which features should be included in disaster apps.
RQ2. What do users discuss in disaster app reviews?
Rationale. We want to explore the perceptions and opinions of disaster app users that are expressed through the reviews of these mobile apps. By employing topic modeling techniques, this RQ extracts and categorizes the primary topics discussed in user reviews. This analysis, covering 28,161 reviews across 44 mobile apps, identified 18 distinct topics.
RQ3. What are the common challenges in using disaster apps as reported by their users?
Rationale. We want to develop deeper understanding of key grievances and dissatisfaction among users of disaster mobile apps. This was analyzed with the ’user ratings’ provided for each app, and via a sentiment analysis performed on the reviews.
Figure 1: Overview of our study approach
3.2 Data Collection
Figure 1 summarizes our research methodology. This section details the steps involved in identifying, gathering, and preparing the data to answer our RQs.
3.2.1 A Dataset of Disaster Mobile Apps
To answer RQ1, we needed a collection of disaster apps. Currently, there is no available dataset of such apps. We adopted the following steps to identify disaster apps.
Step 1: App Collection. This step involved locating apps for disasters. We used three approaches to achieve this. The collection process was conducted between 19 October and 25 October 2023. Applications published outside this timeframe were not included in our study.
Approach 1: Literature-based app collection. We first reviewed existing literature to identify studies that compiled lists of disaster management applications. From the works in [Tan2017, Tan2020Modified], we obtained a combined list of 115 mobile apps related to disaster management. At this stage, duplicate entries were not removed, as the deduplication process was performed later after integrating this list with the apps collected through the other two approaches.
Approach 2: Keywords-based app collection. Due to the limited apps that were available via the literature-based app collection approach, we searched for relevant apps in both the Google Play Store and the App Store in different parts of the world. We identified disaster apps using disaster-specific keywords [Navarro2021], such as ‘flood’, ‘earthquake’, and ‘tornado’. These keywords were used to search in the titles and descriptions of apps in the Google Play Store and Apple Store. This process led to an identification of 54 apps tailored to addressing natural disaster situations.
Approach 3: Country-based app collection. Despite conducting searches with disaster-specific keywords, not all potential apps were discovered. This issue arose because some apps which contain information on natural disasters, do not include explicitly disaster-related terms in their names, tags, or descriptions. For instance, ‘VicEmergency’, an app developed by an Australian state government, provides a variety of warnings, including those for natural disasters. However, its name and tags lack any direct reference to ‘disaster’ or specific disaster types. The description of ‘VicEmergency’ merely states it as the official app for emergency warnings and information in Victoria, with tags limited to ‘News and Emergency’. Consequently, such apps did not appear in search results for disaster-related queries. Additionally, apps developed in non-English speaking countries often have names in their native languages, making them hard to find through English keyword searches, such as the German government’s app ‘NINA - Die Warn-App des BBK’.
To address this challenge we adopted search approach 3 that involved conducting a google search for potential disaster support apps in each country. We used the query “Emergency apps in [Country]” as emergency apps often included functionalities related to natural disasters. We did not use “Disaster apps in [Country]” as a query because this would be overly restrictive and may not yield a complete app list. Some of the disaster apps obtained through this method include ‘Yurekuru Call’ and ‘Safety Tips’. These apps were found in a blog written by Robin Lewis [lewis:2018], which was discovered using the query “Emergency apps in Japan.” This approach was repeated for all 196 countries worldwide [britannica].
With approach 3, we found 73 disaster-related apps developed by government and non-government entities, available on the Google Play Store and the App Store. However, this was much less than what we expected for an overall of 196 countries in the world. While it was understandable that some developing countries may not have an app due to resource limitations, we did not feel that this justified all the missing countries. Therefore, to understand why many countries have not developed specific disaster apps, we exhaustively reviewed every official state government website to identify how they provided disaster information to their citizens. Our research showed that each country had various methods of disseminating disaster information, many often without relying on mobile apps. For instance, Austria utilizes sirens  [^1] https://www.bmi.gv.at/204_english/skkm/warning.aspx, Canada informs their citizens via television and radio  [^2] https://www.alertready.ca, Saudi Arabia sends notifications directly to personal devices  [^3] https://998.gov.sa/English/Pages/Automatic-Early-Warning.aspx, India employs SMS  [^4] https://sachet.ndma.gov.in, and Belgium opts for email  [^5] https://www.be-alert.be/en/how-does-be-alert-work. These are only one of the methods that are used by these countries as most countries prefer to employ multiple methods for disseminating disaster information.
Step 2: Combining List of Apps and Removing the Duplicates. Figure 1 shows that 242 apps were obtained from our three app collection methods. This included 115 apps from the literature review, 54 apps from keyword-based searches, and 73 apps from country-based searches. However, there were overlaps in apps that were collected from some methods, thus we remove the duplicates.
Step 3: Filtering Apps. After removing duplicates we get 190 apps. To refine this collection, we implemented a two-stage filtering process. First, we excluded applications that were not available in either the Google Play Store or the App Store, since the absence of availability prevented us from retrieving user reviews. Second, we eliminated false positives [Chembakottu2023]. False positives were defined as applications that did not include functionalities relevant to natural disasters. Specifically, we excluded applications that did not provide at least one feature listed in Table 1. For instance, some applications were designed to facilitate communication with law enforcement authorities. While these address emergency situations, they fall outside the scope of our study, which is limited to natural disaster contexts. Similarly, applications providing assistance in medical emergencies (e.g., cardiac arrest) were excluded because their focus is on general health crises rather than disaster-specific scenarios. This rationale also applied to generic social media and crowdfunding platforms, which, although they may contain incidental information about disasters, are not explicitly developed for disaster management. To ensure the accuracy of this second filtering stage, we downloaded selected applications to verify their functionalities directly. Following this, we identified 44 applications that incorporated features directly related to natural disaster management (See Table 1). The complete list of applications detailed in Appendix Table A2 and Table LABEL:tbl:list_of_play_store_app.
3.2.2 A Dataset of Disaster App Reviews
Collecting Reviews: To answer RQ2 and RQ3, we needed user reviews of the 44 disaster apps. To gather user reviews of 44 disaster apps from the Google Play Store and the App Store, we utilized an app-store-scraper  [^6] https://github.com/cowboy-bebug/app-store-scraper and google-play-scraper  [^7] https://github.com/JoMingyu/google-play-scraper. These tools require the app’s ID, name, and the specific country of the store. The app’s name is visible upon selection, whereas its ID and the relevant country code are identifiable through its URL. We collected 37,272 user reviews from 44 disaster apps.
Translation: The collected reviews encompassed a diverse linguistic range, particularly for applications originating from non-English-speaking countries. Examples include DEC112 2.0 (German), Zachranka (Czech), and Yurekuru Call (Japanese). Among the 44 apps analyzed, 9 contained reviews in languages other than English, amounting to 2,793 reviews that required translation. To address this, we employed the OPUS-MT Zero-Shot Translation [Tiedemann2023] model to translate non-English reviews into English. We selected OPUS-MT because it represents one of the most comprehensive collections of machine translation models. The OPUS-MT repository comprises more than 2,300 models, supporting 4,560 language directions across 294 languages [Tiedemann2023]. Furthermore, OPUS-MT achieves competitive translation quality, with BLEU scores comparable to large-scale systems such as NLLB [nllbteam2022languageleftbehindscaling]
Exclusions: The next step was to exclude brief reviews comprising fewer than four words. The reason behind this exclusion is short reviews offering minimal commentary such as ‘good’, ‘not accurate’, and ‘works well’. Thus, this type of reviews provide insufficient information for meaningful analysis and can’t be used to generate a topic for RQ2. In addition, reviews without a rating were removed, as they could not be used in the sentiment analysis. After this filtering, the dataset was reduced to 28,161 reviews, providing a solid foundation for the analysis of RQ2 and RQ3.
3.3 Methodology Approach
3.3.1 RQ1: Key features of disaster apps
In this RQ, we examine the features available within disaster support apps, the types of natural disasters they address, and their geographic coverage. Subsequently, we analyze the apps’ popularity and quality based on metrics such as the number of downloads and user ratings and explore their monetization strategies.
Analysis 1- App Features: To identify the available app features, we conducted a hands-on evaluation by downloading and using each application directly. This approach allowed us to explore the full range of functionalities beyond what was described in the app listings or shown in image previews. During this process, we noted each feature observed through actual use and interaction with the app interface. The collected features were then categorized into three main types, aligned with the four stages of the disaster lifecycle: preparation, response, recovery, and mitigation [04Unit4d19_online].
Analysis 2- Disaster Type Coverage: In addition to categorizing apps based on their lifecycle phase, we also examined the types of disasters each app addresses. Apps were classified into two groups: single-hazard apps and multi-hazard apps. Single-hazard apps have features that focus on only one type of disaster, such as an earthquake, while multi-hazard apps cover multiple types of disasters.
Analysis 3- Geographic Coverage: Apps can also be classified based on the geographic scope of their coverage. We categorized the apps into four levels: local, national, regional or territorial, and global. Local apps provide coverage for areas smaller than a nation, such as a state, province, or city. National apps cover a single nation. Regional or territorial apps cover multiple nations or areas, such as the Pacific regions. Global apps provide coverage all nations.
Analysis 4- Popularity and Quality Analysis: The popularity of an app can be gauged by its number of downloads, while user ratings offer insights into its perceived quality. These data can be obtained from the homepage of each app on the Google Play Store and the App Store. For example, Figure 2 shows that the ‘Hurricane Tracker Pro’ app has been downloaded 5K+ and has a rate of 4.7.
Figure 2: App pricing scheme, number of downloads, and rating
Analysis 5- Monetization Strategies: Not all disaster apps are freely available. To determine whether an app is free, we examined the homepage of each app, which typically indicates if the app contains advertising or offers in-app purchases (payments within the app to access certain features or remove advertisements). For example, as shown in Figure 2, users must pay \$2.99 to download the ‘Hurricane Tracker Pro’ app. In contrast, the ‘Wind Map Hurricane Tracker 3D’ app, also shown in Figure 2, is free to download but contains advertising and offers in-app subscriptions.
3.3.2 RQ2: User Discussions in App Reviews
To understand the topics discussed by users of disaster apps, we applied BERTopic [grootendorst2022bertopic] to our review dataset. We then performed cluster merging, as clustering algorithms rely on distance metrics and do not inherently recognize when separate clusters represent the same topic. This allowed us to combine similar clusters. After merging, we refined the clusters by removing outliers within each cluster. The overall topic modeling process is illustrated in Figure 3, and the details of the approach are described below.
Figure 3: Topic Modeling Process
BERTopic: In this research, BERTopic was utilized for topic modeling, as it is a state-of-the-art model for clustering tasks. BERTopic converts documents into vectors using the Sentence-BERT (SBERT) [reimers2019sentencebertsentenceembeddingsusing] framework. SBERT is utilized for document embedding because it can retain semantic information. BERTopic provides two options for sentence embedding: all-MiniLM-L6-v2888https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 for English sentences and paraphrase-multilingual-MiniLM-L12-v2999https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 for multilingual sentences. The first option was chosen because we translated all non-English reviews into English, so we don’t need a multilingual model for sentence embedding. Subsequently, the dimensionality of the embedding vector is reduced using UMAP (Uniform Manifold Approximation and Projection) [mcinnes2020umap].
UMAP was selected for dimension reduction in BERTopic for its superior embedding stability and significantly faster runtime compared to other methods such as T-SNE [JMLR:v9:vandermaaten08a]. Once reduced, the embedding vector is clustered using HDBSCAN (Hierarchical density-based clustering) [mcinnes2017hdbscan]. HDBSCAN was chosen because this model can prevent unrelated documents (outliers) from being assigned to any cluster and is more robust in parameter selection. HDBSCAN has parameters that need to be specified, among which is min_topic_size [grootendorst2023parametertuning]. This parameter ensures that each cluster has at least min_topic_size members. If there are numerous points in close proximity but fewer than the min_topic_size, these points will either join another cluster or become outliers. In this study, the min_topic_size used in this clustering was 50.
Cluster Merging and Naming: After the model training is complete, BERTopic will provide results in the form of topic IDs, topic names, keywords, and the number of members in each topic. To ensure the accuracy of the topic naming, a manual analysis is conducted to rename each topic. This analysis involves examining the keywords of each topic, as well as checking 50 data points that are closest in distance to the center of each cluster. After all clusters are appropriately named, similar clusters are merged, for example, clusters named “app notification” and “app alert” are considered the same topic.
Cluster Refinement: In the topic modeling process, each identified topic is refined by removing reviews that do not align well with the topic’s main theme. This is done by using BERTopic to re-cluster the reviews within a topic into two groups: outliers and non-outliers. Reviews classified as outliers are discarded, ensuring that only the relevant reviews are retained as the final result of the topic modeling process. If a topic contains no outliers, the number of reviews remains unchanged. Finally, every user review is assigned to a single cluster since BERTopic is not soft clustering method. If a review contains multiple topics, we assign it to the most dominant topic.
3.3.3 RQ3: Challenges in Disaster Apps
Figure 4: Sentiment Analysis Process
Sentiment Analysis: This research question examines which topics have the highest percentage of negative reviews. This information can help developers identify dominant user complaints and prioritize their resolution. Ratings are commonly used to indicate user satisfaction with an app. However, Text-Rating Inconsistency (TRI) can occur when users provide ratings carelessly [BinFu2013]. For instance, a user may give a positive rating immediately after downloading the app without actually using it, or a user may submit numerous complaints yet assign a neutral rating, such as a score of three. Relying solely on ratings may therefore fail to accurately reflect user sentiment. TRI can also result in mislabeled data, which reduces the reliability of supervised learning approaches. To address this issue, unsupervised methods that do not depend on label accuracy can be applied, with VADER [hutto2014vader] being a particularly effective option.
The process of sentiment analysis is shown in Figure 4. The outcome of VADER includes four types of sentiment scores: negative scores, neutral scores, positive scores, and compound scores. VADER considers that each sentence may contain both negative and positive elements. For example, consider the review “Currently, I like the app very much, but you should be able to change the language within the app.” This review contains a positive element: “I like the app very much” but also a negative element: “you should be able to change the language”.
VADER provides a combined score of these elements, known as a compound score, in addition to the individual sentiment scores. This compound score, ranging from -1 to 1, will be used as the result of sentiment analysis in this study. We will classify sentiment as positive (score >= 0.05), neutral (0.05 > score > -0.05), or negative (score <= -0.05), following similar approaches used in prior work [Chembakottu2023].
Distribution of review sentiment: Our analysis divides sentiment into three categories: positive, neutral, and negative. We observe the number of reviews for each sentiment and their corresponding percentages.
Topic with Lowest Rating and Sentiment Score: In this analysis, we examine the user rating and sentiment score for each topic. Next, we identify which topics have the lowest score, enabling developers to prioritize these topics for improvement.
4 Results
4.1 RQ1: Key features of disaster apps
Analysis 1 — App Features: Each disaster support app provides one or more features related to natural disasters. The following section explains the key features we categorized according to the stages of a disaster event: pre-disaster (preparation), during-disaster (response), and post-disaster (recovery). The number of apps offering each feature type is shown in Table 1.
Table 1: App feature categorization and number of apps
Category Feature No. of Apps
Preparation Early Warning Alerts 10
Disaster Forecasting 7
Prep Tips/Tutorials 8
Response Real Time Alerts 30
Disaster Maps 21
Real-Time Tracking 6
Emergency Reporting 1
Recovery Disaster Updates 26
News/Bulletins 8
Share Experiences 3
Safety Confirmation 3
  • Refer to caption
    Figure 5: Features of Preparation Stage
  • Features of Preparation Stage: The user has not yet experienced a disaster during this stage, so the features here are geared towards preparation.
    1. 1.
      Early Warning Alerts: This feature enables users to receive immediate information when a disaster is about to occur, allowing them to take precautionary actions to protect themselves. Apps offering this feature are commonly developed in Japan, particularly those related to earthquakes. A good example is the Yukureku Call app (see Figure LABEL:fig_ew).
    2. 2.
      Disaster Forecasting: This feature supports the prediction of potential disasters at specific times and places. Early warning alerts are sent moments before a disaster and are almost certain to occur. In contrast, forecasting provides a longer lead time, giving users more time to prepare. However, its accuracy is lower than that of early warning systems. This feature is common in apps that monitor wind movements, as typhoons and hurricanes can be predicted more reliably than other disasters. An example is the Seastorm Hurricane Tracker app (see Figure LABEL:fig_forecast).
    3. 3.
      Preparation Tips/Tutorials: This includes education and training material to help prepare for disasters. It can include tips and tutorials for self-rescue and recovery steps. Such an example is shown as advice on earthquake response protocols from the ‘Hazard Red Cross’ app (see Figure LABEL:fig_tips).
    Figure 6: Features of Response Stage
  • Features of Response Stage: When a disaster occurs, several features can be utilized.
    1. 1.
      Real Time Alerts: This alert appears as some form of pop-up, informing the user of a disaster happening in the area according to their radius settings or specified area of interest. This feature differs from an early warning alert because notifications arrive during a disaster, not beforehand. For instance, if a user sets a threshold for earthquakes measuring 3 on the Richter scale and above, they will receive a notification when such an event occurs. Such an example is provided in ‘QuakeFeed Earthquake Tracker’ app (see Figure LABEL:fig_at_disaster_1).
    2. 2.
      Disaster Maps: Such a feature aims to provide some form of geographic-based information to users regarding a disaster occurring at a specific location or coverage area, presented in the form of an interactive map. Such an example can be seen in ‘Hazards Near Me NSW’ app in Figure LABEL:fig_at_disaster_2).
    3. 3.
      Real-Time Tracking: This feature provides live updates on ongoing disaster events or wind conditions, as it is available for wind-related disasters such as hurricanes, cyclones, and typhoons. Therefore, the app will display animated representations of wind-related disasters, which typically occur in the ocean. An example can be seen in ‘Wind Map Hurricane Tracker 3D’ app (see Figure LABEL:fig_at_disaster_3).
    4. 4.
      Emergency Reporting: This feature aims to facilitate emergency calls or reporting when a disaster strikes. This feature was only available on ‘SES Assistance QLD’ app in our dataset (see Figure LABEL:fig_at_disaster_4).
    Figure 7: Features of Recovery Stage
  • Features of Recovery Stage: The key features available in the post-disaster stage.
    1. 1.
      Disaster Updates: Information regarding disaster events can be presented in the form of short updates. This usually consist of a list format, such as a list of all the earthquakes in the past month. In such a feature, users can view disaster statistics, such as the magnitude of the disaster, the depth from the ocean, and the distance to the nearest city. A good example is that used in ‘Earthquake Alert!’ app (see Figure LABEL:fig_rec_disaster_1)
    2. 2.
      News/Bulletins : Information regarding disaster events can also be presented as comprehensive news bulletins. These provide post-disaster information with comprehensive analysis and reporting, such as the number of people injured, the count of collapsed buildings, and the economic losses incurred in the areas affected by the earthquake. News bulletins tend to be longer than short disaster updates. One app that provides this feature is ‘Disaster Alert’ app (see Figure LABEL:fig_rec_disaster_2).
    3. 3.
      Share Experiences: This feature can be utilized to notify others when we feel a disaster. For example, in the ‘Volcanoes & Earthquakes’ app, there is an ‘I Felt It’ button that appears when the app launches a disaster notification (see Figure LABEL:fig_rec_disaster_3).
    4. 4.
      Safety Confirmation: Some apps offer safety confirmation features, such as the ‘National Evacuation Center Guide’ app. This app allows users to confirm their safety to friends or family through contact tracing, location sharing, and short messages within a safety confirmation group in the app (see Figure LABEL:fig_rec_disaster_4).
Table 2: Disaster Type Distribution of Mobile Apps (The Google Play Store and The App Store)
Type of Disaster Number of Apps
Multi-Hazard   26
Earthquakes   16
Wind-Related Disaster   8
Floods   3
Fire-Related Disaster   1
Table 3: Geographic Coverage Distribution of Mobile Apps (The Google Play Store and The App Store)
Geographic Coverage Number of Apps
Global   22
Region / Territorial   6
National   20
Local   8
Analysis 2 — Disaster Type Coverage: Based on the types of disasters covered, the apps can be divided into two categories: multi-hazard apps and single-hazard apps. Multi-hazard apps cover various types of disasters. For example, the Disaster Alert app covers 18 types of disasters. In contrast, single-hazard apps focus on only one type of disaster. From our review, we found that some apps focus solely on earthquakes, some on wind-related disasters, others on floods, and some on fire-related disasters. The distribution of app types based on the disasters they cover is shown in Table 2. Detailed information on which disasters each app covers can be found in Table A2 and Table LABEL:tbl:list_of_play_store_app.
Analysis 3 – Geographic Coverage: Geographic locations supported in disaster apps can be categorized as global and non-global. The non-global category includes regional, national, and local apps. Global apps can be accessed from any location and cover all countries. In contrast, non-global apps are limited to specific geographical areas. Regional apps cover broad areas such as the USA and Canada or the USA and Europe (see Appendix Table A1). Some regional apps also extend across oceanic basins, especially those focusing on wind-related disasters. National apps target a single country only. Territorial apps, such as ‘VicEmergency’, focus on a sub-national area like Victoria, Australia, while ‘SD Emergency’ covers only San Diego, California, USA. The distribution of apps based on coverage is shown in Table 3. Detailed lists of these apps and their geographic coverage can be found in Appendix Table LABEL:tbl:list_of_play_store_app for the Google Play Store and Table A2 for the App Store.
Analysis 4 – Popularity and Quality: On the Google Play Store, the rating of an app is given in a numerical range from 1 to 5, and the number of downloads is expressed within specific ranges, such as 10K+, 500K+, 1M+, and 5M+. Meanwhile, on the App Store, the app rating follows the same scale, ranging from 1 to 5. However, the App Store provides precise figures without rounding to the nearest thousand or million for the number of downloads. The specific number of downloads for each of the apps we explored can be found in Appendix Table A2 and Table LABEL:tbl:list_of_play_store_app.
Table 4: Comparison of Total Downloads and App Ratings
Variable Statistic The Google Play Store The App Store
Rating Mean 3.88 3.68
Median 4.10 3.75
Min 2.10 1.70
Max 5.00 4.80
Downloads Mean 559K+ 840
Median 100K+ 435
Min 5K+ 6
Max 5M+ 6,400
Table 5: Pricing Schema of Natural Disaster Apps
Pricing Schema
The Google
Play Store
The
App Store
Free 22 12
Free with Ads 5 0
Free with In-App Purchase 3 3
Free with Ads and In-App Purchase 8 0
Paid 1 1
Paid with In-App Purchase 1 0
Of the 44 apps in this study, some of are published in both the Google Play Store and the App Store. The number of disaster-related apps in the Google Play Store is 40, while in the App Store, there are 16. On average and median (see Table 4), the rating value in the Google Play Store surpasses that of the App Store. The average and median ratings on the Google Play Store are 3.88 and 4.1, respectively, whereas on the App Store, they stand at 3.68 and 3.75. The app, ‘BD999’, with the highest rating, is also on the Google Play Store, which received a perfect rating of 5. Meanwhile, the app, ‘SES Assistance QLD’, with the lowest rating, is on the App Store, which received a rating of 1.7.
The number of downloads for all disaster-related apps in the Google Play Store exceeded 22 million, while in the App Store, it is over 13,000. Statistically, the average and median number of downloads on the Google Play Store are over 559,000 and 100,000, respectively, while on the App Store, they are 840 and 435. What’s interesting here is that the app with the lowest number of downloads on the Google Play Store has almost the same value as the one with the most on the App Store. However, the significant difference in the number of downloads in these two stores is reasonable, considering the higher percentage of Android users compared to iOS, with Android users accounting for 70.11% and iOS users for 29.19% [StatistaMobileOS]. The app ‘NINA - Die Warn-App des BBK’ has the most downloads is in the Google Play Store with more than 5 million downloads. The app ‘SES Assistance QLD’ has the fewest downloads is in the App Store, only downloaded 6 times.
Analysis 5 – Monetization Strategies: Developers can earn revenue from the apps they create through advertisements, in-app purchases, paid downloads, or a combination of these methods. More than half (22 apps) of the disaster apps were free, while the remaining apps adopted various monetization strategies. As shown in Table 5, five apps on the Google Play Store generate income through in-app advertisements. Three apps offer premium feature packages or provide options to remove ads. Eight apps include both advertisements and premium features available for purchase. One app requires payment upon download, while another requires payment and also offers an upgrade option. Meanwhile, the primary monetization approach in the App Store involves offering in-app purchase services, either as free apps with optional purchases or as paid downloads. The detailed monetization strategy for each app is provided in Appendix Table A2 and Table LABEL:tbl:list_of_play_store_app.
4.2 RQ2: User Discussions in App Reviews
To determine the optimal number of clusters, we evaluated using the Dunn Index and Davies-Bouldin Index. A higher Dunn Index indicates that the clusters are more well-separated, whereas a lower Davies-Bouldin Index indicates better clustering results. We tested a range of topics from 10 to 30. After the experiments, the results are shown in Figure 8. From this figure, it can be seen that the optimal number of clusters is 22.
Figure 8: Evaluation Metrics of BERTopic
However, the automatic topic names generated by BERTopic were not always representative, such as ‘interface_features_ui_layout’ and ‘fix_please_issues_problems’. We therefore renamed these clusters as User Interface and Bug Fixes. After all clusters were appropriately named, similar clusters were merged; for example, clusters named “app notification” and “app alert” were considered the same topic.
Following this merging process, we obtained 18 distinct topics. These topics were then ranked according to the number of reviews per topic, as shown in Table 6. In that table, we provide the topic ID, topic name, keywords representing each topic, total number of reviews in the topic, and review length. The review length in the table refers to the median number of words per user review. We use the median instead of the mean because it is more robust for skewed distributions.
Among the identified topics listed in Table 6, two features—the Alert feature and the Watch Zone feature—stand out as unique to disaster-related applications and are not observed in other types of apps. Notably, the alert functionality is the most frequently discussed theme. A detailed discussion of each topic generated by BERTopic, along with representative examples, is presented in the following section.
Table 6: Topic Extracted from User Reviews
Topic ID Topic Keywords Total Review Length
1 Alert Functionality alert, notification, warning, report 3,129 17
2 Map Usage map, satellite, view, landscape 567 16
3 Monetary Transactions and Ads money, cost, refund, ads, price 542 20
4 Location Functionality location, GPS, enabled 536 22
5 Disaster Updates know, informed, world 499 22
6 App Crashes crash, wont, open, closes 491 17
7 Alert Sound beep, sounds, loud, alarm, siren 462 20
8 Slow Performance heavy, slow, unresponsive, lag, sluggish 363 21
9 Bug Fixes please, fix, repair, problem 285 25
10 App Configuration configuration, settings, change, setup 246 17
11 User Interface hard, see, font, small, interface 227 18
12 User Safety safe, security, family, stay 214 22
13 Sign Up and Sign In Problems register, failed, password, email 155 22
14 Network Connectivity network, error, connect, server 140 20
15 Watch Zone Functionality watch, zone, warnings, notified 139 27
16 Battery Drainage battery, consumption, drain, energy 136 19
17 Reliability reliability, inaccurate, inconsistent 105 18
18 Language Preferences translation, German, English, speak 70 21
4.2.1 Identified Topics
1) Alert Functionality: The most common topic discussed by users is the alert functionality. This includes 3,129 reviews (see Table 6). Users mainly express their satisfaction with the alert functionality because they are informed quickly during disasters, as exemplified in the review “I absolutely love this app. It’s accurate and informative. I check it daily for New Zealand earthquake and volcanic activity, and I receive alerts/notifications. News comes through almost instantly when earthquakes occur”
However, there are some reviews in which users were disappointed due to the late delivery of alerts. An example of such a comment is “… I didn’t receive timely warnings. Alerts can be sent between 10 minutes to an hour late after an earthquake occurs!”
Another common complaint is some users not receiving alerts when disasters occur. One review related to this issue is “… This week has been full of severe thunderstorms, flood warnings, and tornado warnings, yet I have not received any notifications or alerts from this app…”
2) Map Usage: The second most-discussed topic concerns map usage, comprising 567 reviews. Common complaints under this topic include insufficient map detail, the clustering of alerts during zooming, and various issues related to functionality and usability. For instance, one user commented, “… The UI is incredibly clunky, maps are extremely slow to load, and warning icons disappear when moving across the map…”
However, some users were satisfied with the map feature in the app. They are happy about aspects such as accuracy of the maps, fastness in loading, and ease of reading and usage. This is shown in the review “… The map feature works very well, and zooming in provides great detail. Positioning the map for a particular global view also works very well…”
3) Monetary Transactions and Ads: This topic contains customer reviews regarding subscriptions, payments, and advertising. Some users give positive ratings and comments because the app is free and does not have ads, such as “Brilliant. This is actual free app… It’s the first 5-star app for me… Plus, there are no ads!”
However, many other users complain about the disruptive ads, for example “… with the latest update, you don’t just get banner ads,but you also get forced full-page pop-up ads, which are incredibly obtrusive and annoying.” Other users also expressed dissatisfaction after purchasing the app or upgrading to the Pro version, stating that the additional features did not meet their expectations. For example, one review noted, “I paid for the Pro version, and there’s nothing special about it. The app doesn’t warn you right away; it has like a 10–15 minute wait time before it pops up. Living in Alaska, this app is worthless to me.”
4) Location Functionality: This topic contains user reviews related to location functionality, such as satisfaction regarding the ability to add their preferred location to be monitored. This is shown in review “I have friends and family scattered all over the country. The app allows me to add the places where they live and provides alerts if there are hazards such as wildfire weather, earthquakes, or hurricanes approaching. It’s easy to customize the alerts I want and to add new places. It’s also great that it will monitor wherever I am.”
Several other reviews mention user dissatisfaction due to inability to add locations or inability to find locations. One such review is “The app will not allow me to add a home location or turn on live monitoring, rendering the entire app useless. My location permissions are set to ’always’, and the app can detect my location via GPS. However, saving never works and instead always produces an error message.”
5) Disaster Updates: This topic includes reviews from users who are mainly pleased to stay informed about disasters occurring in various parts of the world. Some examples are “Great to be able to look all over the world and see what is happening!”, “…The information it contains offers a full view of what earth is doing. Checking out this app is like watching the earth breath and grow or gag and stumble…”. and “… Since I have friends all around the world, I like knowing where the earth is quaking. At times, it is shocking to know how many tremors and earthquakes happen everyday, but I would rather know, than not.”
6) App Crashes: This topic contains user reviews related to the app being unusable due to crashes. Almost all reviews are negative. App crashes can occur in different stages of using the app. One instance is when users attempt to open the application “… lately, either it won’t open at all, or it closes after a couple of seconds…”. It can also happen in other stages of using the app such as “… Sometimes it crashes when I am in the settings, and I have to restart it…”
7) Alert Sound. This topic discusses issues in the level of alert sounds. Sometimes the issue is with the loudness of the alert sound “… but the notification sound is loud and annoying. I turned down my notification volume on my phone, which made other apps quieter, not what I wanted. However, within 2 days, the volume for this app was so high-pitched that it hurts your ears, and the volume on other apps was still low.”. Some users had complaints about the absence of sound, causing users to be unaware if a disaster occurs “A tornado alert was issued by this app, but there are no sounds at all! A wind chime would be more useful than this app. An audible alarm that alerts you should be mandatory on all weather apps.”
8) Slow Performance: This topic contains user reviews related to slow app performance experiences. It can involve slow loading features or slowness when using the features. Example reviews include “Worst app. If you want to view the map and look at details, you have to wait 10 minutes. Super slow.” and “… it’s very slow trying to move around the app, especially when trying to make a watch list…”
9) Bug Fixes: This topic contains customer reviews regarding bug fixing, including both request to fix bugs and appreciation about bugs that have been resolved. Example of a request to fix bugs is “I have not received any notifications from this app for months! I have logged out, uninstalled, and reinstalled multiple times, but still nothing… Please fix it!”. Example of an appreciation about a bug fix is “Thank you very much for fixing the app! It’s the best one for looking up earthquakes and has a great interface and filters!”
10) App Configuration: This topic contains user reviews related to app configuration. Users may express satisfaction with the flexible configuration, or they may encounter problems with the application configuration. Satisfied customer review can look like this “… a fantastic tool. It offers many settings for configurations tailored to specific needs.” Problems in configuration can be “It won’t open settings when I click, and I can’t adjust which warnings or emergencies I want alerts for.”
11) User Interface: This topic contains user reviews related to the app’s interface. Many users may express satisfaction with the UI such as “Great clean new look. Looking forward to future updates, more new features, and improved design.”. They also highlight problematic issues in UI aspects, such as color and size. One review particularly from a user with vision issues stated “The color combination of black with dark blue is terrible. I would have given it five stars if I could actually read some of the menu labels at the bottom of the screen. My eyesight is limited to one eye, with about 65 to 70% vision out of that eye. I need good contrast to be able to see clearly.”
12) User Safety: This topic contains reviews from users who are happy because they feel safe and helped by this application. An example reviews are “… puts a lot of safety ideas for different emergencies in one place. I really like the monitoring feature, with family scattered or traveling a lot” and “Last night we had an earthquake 40km away had this app on my phone unopened for many months but still got a notification seconds before we felt it. It allowed us to immediately recognise what was happening and start heading to a safe area. Quite impressed!”
13) Sign Up and Sign In Problems: This topic contains user reviews complaining about not being able to register or problems during the login process, with almost all reviews negative. Examples are “Keeps saying registration failed… Already missed a fire near me due to this new app. Please fix.” and “It won’t even let me log into the app at all. It keeps saying my login credentials are incorrect, but when I use the same login info on the web browser, it logs right in. I uninstalled it and won’t be reinstalling it!!”
14) Network Connectivity: This topic contains user complaints regarding the inability to use the app due to connectivity issues, almost all reviews negative. For example, as shown in “Keeps telling me no internet connection detected. Yet I can do everything else and use all other apps without any problems.” Another example of a problem is a user failing to connect to the app server “Many times I get an error message about not being able to connect to the server…”. Some users requested the offline mode for the app as shown in review “Great app, although it’s pretty limited without power or internet connectivity. It would be better if it could be mirrored locally for offline use.”
15) Watch Zone Functionality: This topic contains the longest reviews compared to other topics, with the median word count being 27 (see Table 6). This topic contains complaints about the ‘watch zone’ not being able to be set such as “Almost useless. You cannot create watch zones. It just says ’unable to create watch zones’ over and over…”. Besides that, this topic also contains complaints about the watch zone not working properly as shown in review “It gives me notifications every time there’s a thunderstorm anywhere in the state, but when there was actually a fire in my watch zone, it didn’t come up.”
16) Battery Drainage: This topic is quite straightforward, covering the issue of the app consuming a lot of battery, with almost all reviews highly negative. “… the app was using a huge amount of battery and data in the background. Thirty-one percent battery drain was attributed to this app, along with 337MB of data in only a few days of background activity.”
17) Reliability: This topic contains customer reviews regarding whether users find the app reliable or not. Example positive perceptions look like e.g. “The app is very reliable and trustworthy. I depend on it as I live in a quake-prone area, and the information that this app provides is very important to me and my family.” A negative perception is reflected in the review “Unreliable for personal safety and potentially life-threatening during the recent 2020 fires for many people. I don’t know if the networks were jammed or if the app was broken, but we had to rely on looking out a window for a true assessment of whether there are fires near me or not.”
18) Language Preferences: This topic contains customer reviews related to language issues. Many users wish for the app to be available in a language they understand or are fluent, as reflected in the review “… The only criticism I see right now is the lack of supported languages. I think an English version should at least be provided. Other languages could be added by volunteers, for example, through open-source contributions or via a "donate a translation" button in the app.” Some users suggested alternatives for in app translation : “Could you make it possible to have the English translation? Or at least make it possible to copy the text so that I can translate it.”
4.3 RQ3: Challenges in Disaster Apps
Table 7: Sentiment Analysis by User Rating and Sentiment Score
Number of Reviews [%]
User Rating
Negative   2663 [24.30%]
Neutral   1911 [17.44%]
Positive   3732 [44.93%]
Sentiment Score
Negative   2663 [24.30%]
Neutral   1233 [11.25%]
Positive   4410 [53.09%]
Distribution of Review Sentiment: Table 7 presents the distribution of sentiment analysis results across the user review dataset. The proportion of positive sentiment derived from user ratings is 44.93%, while the proportion based on sentiment scores is slightly higher at 53.09%. The percentage of neutral sentiment in user ratings is 17.44%, which is marginally higher than that from sentiment scores at 11.25%. Both methods indicate an identical proportion of negative sentiment at 24.30%.
Table 8: Average rating, sentiment score, and combined sentiment–rating score for each topic.
ID Topic Avg. Sentiment Score Avg. User Rating Avg. Sentiment- User Rating
1 Alert Functionality 3.47 3.48 3.48
2 Map Usage 3.64 3.48 3.56
3 Monetary Transactions and Ads 3.40 3.01 3.21
4 Location Functionality 3.34 2.86 3.10
5 Keep Updates 3.86 4.66 4.26
6 App Crashes 3.24 2.12 2.68
7 Alert Sound 3.33 3.47 3.40
8 Slow Performance 3.27 2.86 3.07
9 Bug Fixes 3.59 2.85 3.22
10 App Configuration 3.18 2.89 3.04
11 User Interface 3.75 3.81 3.78
12 User Safety 3.55 3.71 3.63
13 Sign Up and Sign In Problems 2.94 1.65 2.30
14 Network Connectivity 2.99 2.28 2.64
15 Watch Zone Functionality 3.23 2.53 2.88
16 Battery Drainage 3.25 2.47 2.86
17 Reliability 3.33 3.17 3.25
18 Language Preferences 3.37 2.89 3.13
Topics with Lowest Rating and Sentiment Score: Table 8 presents the average sentiment scores and average user ratings for each topic. The user rating ranges from 1 to 5, while the sentiment score originally ranges from -1 to 1. To make them comparable, the sentiment scores were rescaled to a 1–5 scale using the formula $(x+1)\times 2$. We also added a column called Avg. Sentiment–User Rating, which represents the mean of the sentiment score and user rating. This helps identify topics that users perceive most negatively based on both sentiment and rating. The topic with the lowest combined score is Sign Up and Sign In Problems, with an average score of 2.30. The second lowest is Network Connectivity with 2.64, followed by App Crashes with 2.68. Two other topics fall below the neutral threshold of 3.00, namely Battery Drainage (2.86) and Watch Zone Functionality (2.88). These results suggest that developers aiming to improve disaster applications should prioritize addressing these issues to enhance user satisfaction.
5 Discussions and Recommendations
Expanding the Scope of Disaster Management Features in Mobile Apps: In disaster management applications, the focus often lies on functionalities that are active during or immediately after a disaster, such as real-time alerts, disaster maps, and updates. While these are undoubtedly crucial, there is a significant opportunity to expand the feature set to better prepare and support users throughout the entire disaster lifecycle. For example, early warning alerts, which are present in only 22% of the surveyed applications, can provide critical advance notice to users, allowing them to take necessary precautions and potentially save lives.
Moreover, the recovery and mitigation phases are equally important, yet often underrepresented in current applications. Features like shelter information, safety confirmations, and recovery actions are vital for helping individuals locate assistance, confirm the safety of friends and family, and understand the next steps they can take. Enhancing the availability of these features could significantly improve the efficacy of disaster response and recovery efforts, addressing gaps that currently leave users under prepared or insufficiently supported after a disaster strikes. By integrating a more comprehensive set of features that span the pre- and post-disaster phases, developers can create more robust applications that cater to a wider range of user needs and scenarios.
Internationalization: As shown in Table 2, among the various apps categorized based on their geographic coverage—22 global apps, 6 region/territorial apps, 20 national apps, and 8 local apps—only the global and regional apps consistently support the English language. Among the national apps from non-English speaking countries, such as ‘Anhaar’ from Mongolia, ‘Emergency Ready’ App from South Korea, ‘BD 999’ from Bangladesh, ‘Alertswiss’ from Switzerland, ‘NINA - Die Warn-App des BBK’ from Germany, and several from Japan including ‘NERV Disaster Prevention’, ‘Safety Tips’, ‘National Evacuation Center Guide’, and ‘Yurekuru Call’, only ‘Safety Tips’, ‘NERV Disaster Prevention’, and ‘Emergency Ready’ App offer English language support.
This limited availability poses significant usability challenges for non-native speakers. For example, some users have expressed frustration with the German app ‘NINA - Die Warn-App des BBK’ for not offering an English version, highlighting the app’s importance and the expectation for broader language support. Similar concerns have been raised about ‘Alertswiss’, where some users have noted the absence of support for all official languages and limitations in using external tools for translation. Another point of contention is the quality of translations, with some users describing them as inadequate.
These examples underscore the importance of incorporating at least one universally understood language, such as English, in disaster management apps. This feature is essential for ensuring the usability of the app especially for visitors and temporary residents. Providing quality translations or, alternatively, enabling features that assist with language translation, like text copying, can significantly enhance the functionality and user experience of these apps.
Monetization Techniques and their Impact: Disaster apps play a crucial role in preparing communities for emergencies and enabling timely evacuations. However, the implementation of certain pricing schemes or the presence of advertisements can hinder their effectiveness. In particular, displaying ads during an ongoing disaster can significantly impact the user experience. Ads can be distracting or cause additional stress, potentially leading to delays in accessing crucial information, which in turn, can jeopardize user safety. It is advisable for developers to consider disabling ads during critical situations to ensure that users can react quickly.
Regarding app monetization, from the user reviews, it is known that while some users are willing to endure few ads for a free service, others prefer a paid model without getting ads at all. To address these divergent preferences and needs, developers should consider flexible pricing models. One effective strategy could be offering a basic free version with non-intrusive ads and an optional subscription-based model that removes ads and offers enhanced features. This approach allows users to choose the version that best fits their needs and financial situation. Moreover, developers should ensure that essential disaster-related functionalities remain accessible to all users, regardless of the version they are using. By providing options for customization and control, developers can create a more user-friendly app that accommodates diverse user preferences while still generating revenue.
Addressing Topics with the Highest Percentage of Negative Reviews: Based on user ratings and sentiment scores, the topics with the lowest values are Sign Up and Sign In Problems, Network Connectivity, App Crashes, Battery Drainage, and Watch Zone Functionality. Except for Battery Drainage and Watch Zone Functionality, the other three topics represent critical usability issues, as they prevent users from accessing or properly operating the app. These problems should be thoroughly resolved during the quality assurance (QA) testing stage to ensure a stable and functional user experience. Regarding Battery Drainage, developers could optimize background processes or implement efficient notification mechanisms to reduce power consumption. For Watch Zone Functionality, developers should allow users to customize or select their preferred watch zones, as this flexibility would improve the app’s relevance and usability across different geographic regions.
Better Support for Diverse App End Users and Platforms: Several user reviews highlighted inadequate support for diverse users of disaster apps. First, in terms of language, the app should provide a global language option, such as English, to accommodate travelers and foreign residents in different countries. Regarding accessibility, some users reported difficulties due to color vision impairments, as they were unable to distinguish certain colors on the map. In addition, small font sizes made it difficult for users with myopia to read the text. For audio alerts, users mentioned that excessively loud notifications disturbed individuals with sensitive hearing, while overly quiet alerts were ineffective for users with hearing impairments. Furthermore, some users reported compatibility issues with specific operating systems, preventing them from using the app. Lastly, many disaster apps require a stable or high-speed internet connection, which can limit access for users in rural or low-connectivity areas. Addressing these issues would make disaster apps more inclusive and accessible to a wider range of users.
6 Threats to Validity
6.1 Internal Validity
App Filtering in Dataset: Constructing the dataset of disaster apps posed specific challenges that could affect the validity of our study. The filtering process risked excluding relevant apps with disaster-related features due to misclassification. To address this issue, we downloaded and tested each app directly to verify its relevance and functionality. Furthermore, all three authors engaged in regular discussions to refine the inclusion criteria, ensuring that the final dataset accurately represents disaster management applications.
6.2 Construct Validity
Methods: In this study, our analysis relied on a single clustering technique, BERTopic, and one method for sentiment analysis, a rule-based model called VADER. Relying solely on these methods might limit the generalizability of our findings, as different techniques could potentially yield varying results.
Model Configuration: The configuration and fine-tuning of BERTopic might have also influenced the results. The hyper-parameters such as the sentence embedding method, distance metric formula, and minimum topic size can significantly affect what the model identifies as a topic. Similarly, the configuration of the VADER model, particularly the sentiment thresholds set for classifying texts as positive (score >= 0.05), neutral (0.05 > score > -0.05), or negative (score <= -0.05), might impact the sentiment analysis. Since VADER is rule-based, there are no hyper-parameters to tune, but the thresholds chosen are crucial in determining the sentiment classification.
These factors suggest a need for caution in interpreting the results. Future studies could benefit from incorporating multiple clustering and sentiment analysis techniques to validate the findings. For sentiment analysis, comparisons between rule-based methods like VADER and machine learning techniques could be particularly illuminating. However, the application of machine learning techniques would require access to well-labeled data, which is essential for training and testing the models accurately.
6.3 External Validity
App Dataset: Our compilation of disaster-related apps utilized literature-based, keyword-based, and country-based methods. Despite these comprehensive strategies, some relevant apps may not have been included due to inaccessible literature, incomplete keywords, or inadequate country-specific queries. These limitations might affect the comprehensiveness and generalizability of our findings, potentially omitting key apps used globally for disaster management and response.
Review Scraping: Data scraping was performed using both the Google Play Store and the App Store scrapers, employing batch scraping strategies to manage access limitations imposed by Google and Apple. We paused for 5 seconds after every 500 reviews, but despite these precautions, some reviews may still have been missed if access was denied due to scraping limits.
Translation: Due to limited resources, our translation efforts focus solely on reviews from apps developed in non-English-speaking countries. While we acknowledge that apps from English-speaking regions may also contain reviews in various languages, we do not translate these in this study. This decision may affect the depth and breadth of our understanding of user feedback from different linguistic backgrounds, although the number of non-English reviews is relatively small.
6.4 Reliability
To enhance the reliability of our study and ensure reproducibility for future researchers, we have compiled a comprehensive replication package openly available on GitHub 101010github.com/Syukrondzeko/natural-disaster-app-review-analysis, which includes the dataset used in our analysis and the source code of clustering and sentiment analysis models utilized. This repository is designed to facilitate replication of our study, providing step-by-step guidelines on how to use it, thus minimizing the risk of obtaining different results.
7 Conclusion
Disaster apps represent a group of apps that are critical for human survival. In this study we conducted a detailed analysis of 44 apps along with their reviews. The analysis of the app features resulted in creating a list of features allocated to three stages of disaster lifecycle: preparation, response, and recovery. This pool of features will help emergency agencies and future app developers to get an overview of the features that can be included in disaster apps. The topics we have identified in RQ2 and the users reactions to these features from RQ3, will help future developers understand common issues that exist in disaster apps and take precautionary steps to prevent them in their apps. These RQ results also show which features were most preferred by users and by incorporating these features the future app developers can increase their user satisfaction.
8 Acknowledgment
Syukron, Madugalla and Grundy were supported by ARC Laureate Fellowship FL190100035.
APPENDIX
Table A1: Region-Specific Disaster Apps
App Name Land Area Oceanic and Basin
USA
Canada
Europe
Atlantic
Ocean
Indian
Ocean
Pacific
Ocean
Central Pacific
Basin
Eastern Pacific
Basin
CodeRED Mobile Alert x x
FloodAlert Waterlevel Alerts x x
Global Storms x x x
Hurricane Tracker x x
SeaStorm Hurricane Tracker x x x
Tropical Hurricane Tracker x x x
Table A2: Overview of Disaster-Related Mobile Apps from App Store
No. App Name Downloads Rating Pricing Geographic Coverage Disaster Type
1 Alert2Me - Emergency Alerts 57 3.2 Free Global Multi-Disaster
2 Alert SA 129 2.3 Free Local Multi-Disaster
3 Alertswiss 522 3.7 Free National Multi-Disaster
4 Disaster Alert 26 3.8 Free Global Multi-Disaster
5 Earthquake 1.7K 4.8 Free with In-App Purchase Global Earthquake
6 Earthquake + Alerts Map & Info 790 4.6 Free with In-App Purchase Global Earthquake
7 Emergency Ready App 15 2.8 Free National Multi-Disaster
8 GeoNet 465 4.8 Free National Earthquake
9 Hazards Near Me NSW 1.1K+ 3.3 Free Local Multi-Disaster
10 LastQuake 104 4.7 Free Global Earthquake
11 My Earthquake Alerts 405 4.6 Free Global Earthquake
12 QuakeFeed Earthquake Tracker 1.1K 4.7 Free with In-App Purchase Global Earthquake
13 Safety tips 13 2.3 Free National Earthquake
14 SES Assistance QLD 6 1.7 Free National Flood
15 VIC Fires 6.4K 4.7 Paid National Fire
16 VicEmergency 614 2.9 Free National Multi-Disaster
Table A3: Overview of Disaster-Related Mobile Apps from Play Store
No. App Name Downloads Rating Pricing Geographic Coverage Disaster Type
1 Alert SA 100K+ 3.5 Free Local Multi-Disaster
2 Alert2Me - Emergency Alerts 10K+ 3.5 Free Global Multi-Disaster
3 Alertswiss 500K+ 4.1 Free National Multi-Disaster
4 Anhaar 10K+ 4.7 Free National Multi-Disaster
5 BD 999 10K+ 5.0 Free National Multi-Disaster
6 CodeRED Mobile Alert 500K+ 2.8 Free Regional Multi-Disaster
7 Disaster Alert 500K+ 4.1 Free Global Multi-Disaster
8 Earthquake 100K+ 4.1 Free with Ads Global Earthquake
9 Earthquake Alert! 1M+ 4.6 Free Global Earthquake
10 Earthquake + Alerts Map & Info 100K+ 4.1 Free with Ads and In-App Purchase Global Earthquake
11 Earthquakes Today 100K+ 4.4 Free with Ads and In-App Purchase Global Earthquake
12 Earthquakes Tracker 500K+ 4.5 Free with Ads and In-App Purchase Global Earthquake
13 Emergency 10K+ 4.2 Free National Multi-Disaster
14 Emergency Ready App 10K+ 3.5 Free National Multi-Disaster
15 Emergency: Severe Weather App 100K+ 3.7 Free National Multi-Disaster
16 FEMA 1M+ 3.8 Free National Multi-Disaster
17 FloodAlert Waterlevel Alerts 100K+ 4.2 Free with In-App Purchase Regional Flood
18 Floods Near Me NSW 50K+ 2.1 Free Local Flood
19 Global Storms 1M+ 4.4 Free with Ads and In-App Purchase Regional Wind-Related Disaster
20 Hazards Near Me NSW 1M+ 2.5 Free Local Multi-Disaster
21 Hazards Red Cross 50K+ 3.1 Free National Multi-Disaster
22 Hurricane & Typhoon Track 100K+ 3.8 Free with Ads Global Wind-Related Disaster
23 Hurricane Tracker 10K+ 4.3 Paid with In-App Purchase Regional Wind-Related Disaster
24 My Alert 50K+ 3.5 Free Global Multi-Disaster
25 My Earthquake Alerts 5M+ 4.7 Free with Ads Global Earthquake
26 My Hurricane Tracker 500K+ 4.7 Free with Ads Global Wind-Related Disaster
27 My Hurricane Tracker Pro 5K+ 4.7 Paid Global Wind-Related Disaster
28 NINA - Die Warn-App des BBK 5M+ 2.6 Free National Multi-Disaster
29 NERV Disaster Prevention 1M+ 4.6 Free with In-App Purchase National Multi-Disaster
30 Natural Disaster Monitor 50K+ 3.5 Free with Ads Global Multi-Disaster
31 National evacuation center guide 100K+ 3.3 Free National Multi-Disaster
32 SD Emergency 50K+ 3.5 Free Local Multi-Disaster
33 SES Assistance QLD 10K+ 3.1 Free Local Flood
34 Safety tips 100K+ 2.3 Free National Earthquake
35 SeaStorm Hurricane Tracker 10K+ 4.1 Free Regional Wind-Related Disaster
36 Tropical Hurricane Tracker 100K+ 4.9 Free with Ads and In-App Purchase Regional Wind-Related Disaster
37 VicEmergency 1M+ 3.3 Free Local Multi-Disaster
38 Volcanoes & Earthquakes 1M+ 4.4 Free with Ads and In-App Purchase Global Earthquake
39 Wind Map Hurricane Tracker 3D 1M+ 4.3 Free with Ads and In-App Purchase Global Wind-Related Disaster
40 Yurekuru Call 1M+ 4.4 Free with Ads and In-App Purchase National Earthquake